A newly issued 10-year maturity, 6% coupon bond making annual coupon payments is sold to the public at a price of $976. What will be an investor’s taxable income from the bond over the coming year? The bond will not be sold at the end of the year. The bond is treated as an original issue discount bond. (Round your answer to 2 decimal places.)
Answer #1

Considering Face value = 1000

Coupon Amount = (6/100)*1000 = 60

ytm = r

PV = 976

976 = 60/(1+r)^1 + 60/(1+r)^2 + 60/(1+r)^3 +60/(1+r)^4 + 60/(1+r)^5 + 60/(1+r)^6 + 60/(1+r)^7 + 60/(1+r)^8 + 60/(1+r)^9 + 1060/(1+r)^10

Solve for r,

r = 6.3312%

B C D F 6.3312% 60 OnMino 60 56.42746 =B4/(1+$D$2)^A3 60 53.06765 =B5/(1+$D$2)^A4 Discounting values to first year 60 49.9078

After one year, 9 years will be remaining in maturity and

The price of the bond will be = 977.794

Taxable income = Price of bond after one year + coupon amount - price of the bond = 977.794 + 60 - 976

= $61.794 Ans